OpenStudy (anonymous):

Jerome currently has an account balance of $1,859.55. His initial deposit on the account was $825 and it earned 5.7% simple interest. How long has Jerome held the account? 18 years 10 years 22 years 40 years

OpenStudy (johnweldon1993):

Well we know the formula for simple interest is I = PRT I = interest earned P = initial balance R = rate T = time We want to solve for time here....so rearrange for T \[T = \frac{ I }{ PR }\] We have P = 825 R = .057 And I = current balance - initial balance = 1859.55 - 825 = 1034.55 So plugging all those in we have \[T = \frac{ 1034.55 }{ 825 \times .057 }\] What does T equal ...?

OpenStudy (anonymous):

AHHH you're the best, thank you so much

OpenStudy (johnweldon1993):

Anytime! :)